OpenStudy (anonymous):

what is the missing term? x/x + 1 , 1 , x+2/x+1, ?

OpenStudy (unklerhaukus):

\[\frac{x}{x + 1} ,\qquad 1 , \qquad \frac{x+2}{x+1},\qquad ?\] have i got these right?

Parth (parthkohli):

The numerator is increasing by 1 every time. Guess now.

OpenStudy (unklerhaukus):

path is using this \[1=\frac{x+1}{x+1}\]

Parth (parthkohli):

A formal way to say this is that the common difference between An and An+1 is 1/x+1.

OpenStudy (anonymous):

x+3/x+1

Parth (parthkohli):

Oops....I guess an answer wasn't needed here.

OpenStudy (anonymous):

i just wanna pinch those cheeks!!!

Parth (parthkohli):

x+1/x+1 when x \(\ne\) -1

OpenStudy (unklerhaukus):

if x=-1 none of the terms would make sense

Parth (parthkohli):

Exactly.